# Analytical bounds for decoy-state quantum key distribution with discrete phase randomization

### v1

@@ III-A BB84 with DPR: Rate analysis
The key rate of the BB84 protocol in Section II-B is lower bounded by [gottesman2004security]
$R^{\text{DPR}}_{\text{BB84}}$ $=\sum_{k=\{0,1\}}Q_{x,k}^{\mu}[1-h_{2}(e_{x,k}^{p,\mu})]-fQ_{x}^{\mu}h_{2}(E_{x}^{\mu}),$ (11)
where $Q_{x,k}^{\mu}$ denotes the gain when Alice encodes the state $|\lambda_{k}\rangle$ with intensity $\mu$, $f$ is the error correction inefficiency, and $h_{2}(x)=-x\log_{2}(x)-(1-x)\log_{2}(1-x)$ represents the binary Shannon entropy. The term $e_{x,k}^{p,\mu}$ denotes the phase error rate for $|\lambda_{k}\rangle$ with intensity $\mu$ as the input state, which in the $X$ basis corresponds to errors caused by Pauli operator $\sigma_{X}$. The second term in the key rate expression accounts for the cost of error correction, where $Q_{x}^{\mu}$ and $E_{x}^{\mu}$ are the total gain and quantum bit error rate (QBER), respectively, for intensity $\mu$. All above terms are for the $X$ basis, which is the main basis used in the efficient QKD protocol we defined in Section II-B.
To compute the key rate, two critical parameters must be bounded: The yield term $Y_{x,k}^{\mu}=Q_{x,k}^{\mu}/p_{k}^{\mu}$, and the phase error rate $e^{p,\mu}_{x,k}$. Both parameters are associated with the quantum state $|\lambda_{k}\rangle$, rather than the photon number state $|k\rangle$ in CPR. Another distinct feature in DPR is that, unlike the CPR case, the yield, gain, and error parameters corresponding to $|\lambda_{k}\rangle$ are now functions of the employed intensity as well. This will make the optimization problem more tedious to solve as we see below.
@@
where the subscript $\gamma=\{x,y\}$ helps us distinguish between gain, yield, and error terms in basis $X$ and $Y$, $e_{\gamma,k}^{b,\alpha}$ is the bit error rate when $|\lambda_{k}\rangle$, with intensity $\alpha$, is at the encoder input.
To minimize Eq. 11 subject to Eq. 12, we use the following inequalities [gottesman2004security, cao2015discrete]:
@@
where $\alpha_{1}$ and $\alpha_{2}$ are two intensity variables in $\{0,\nu,\mu\}$, and the fidelity $F_{\alpha_{1}\alpha_{2}}$ is defined as
@@
Finally, to bound the privacy amplification term, the phase error rate term $e^{p,\mu}_{x,k}$ needs to be upper bounded. We use a virtual protocol described in [gottesman2004security, koashi2005simple] to establish a relationship between the bit error rate $e^{b,\mu}_{y,k}$ in the $Y$ basis and the phase error rate $e^{p,\mu}_{x,k}$ in the $X$ basis, which characterizes the effective basis dependence in realistic scenarios by the following parameter:
$\begin{split}\Delta_{k}^{\mu}=\text{min}(1,\frac{1-F_{k,\text{BB84}}}{2Y_{x,k}^{\mu}}).\end{split}$ (15)
Using a bound proved in [lo2006security], we can get a relationship between $e^{b,\mu}_{y,k}$ and $e^{p,\mu}_{x,k}$ as follows:
$e^{p,\mu}_{x,k}$ $\leq e^{b,\mu}_{y,k}+4\Delta^{\mu}_{k}(1-\Delta^{\mu}_{k})(1-2e^{b,\mu}_{y,k})$
$\quad+4(1-2\Delta^{\mu}_{k})\sqrt{\Delta^{\mu}_{k}(1-\Delta^{\mu}_{k})e^{b,\mu}_{y,k}(1-e^{b,\mu}_{y,k})}.$ (16)
@@
Box 2: Nonlinear optimization problem for DPR MDI QKD with vacuum+weak decoy states Generic Formulation Minimize $R^{\text{DPR}}_{\text{MDI}}=\sum_{k=\{0,1\}}Q_{x,kk}^{\mu\mu}[1-h_{2}(e_{x,kk}^{p,\mu\mu})]-fQ_{x}^{\mu\mu}h_{2}(E_{x}^{\mu\mu}).$ subject to $Q_{\gamma}^{\alpha\beta}=\sum_{k,l=0}^{D-1}p_{k}^{\alpha}p_{l}^{\beta}Y_{\gamma,kl}^{\alpha\beta},$ $Q_{\gamma}^{\alpha\beta}E_{\gamma}^{\alpha\beta}=\sum_{k,l=0}^{D-1}p_{k}^{\alpha}p_{l}^{\beta}e_{\gamma,kl}^{b,\alpha\beta}Y_{\gamma,kl}^{\alpha\beta},$ $|Y_{\gamma,kl}^{\alpha_{1}\beta_{1}}-Y_{\gamma,kl}^{\alpha_{2}\beta_{2}}|\leq\epsilon_{\alpha_{1}\alpha_{2}}+\epsilon_{\beta_{1}\beta_{2}},$ $|e_{\gamma,kl}^{b,\alpha_{1}\beta_{1}}Y_{\gamma,kl}^{\alpha_{1}\beta_{1}}-e_{\gamma,kl}^{b,\alpha_{2}\beta_{2}}Y_{\gamma,kl}^{\alpha_{2}\beta_{2}}|\leq\epsilon_{\alpha_{1}\alpha_{2}}+\epsilon_{\beta_{1}\beta_{2}},$ $e^{p,\mu\mu}_{x,kk}\leq e^{b,\mu\mu}_{y,kk}+4\Delta^{\mu\mu}_{kk}(1-\Delta^{\mu\mu}_{kk})(1-2e^{b,\mu\mu}_{y,kk})$ $\quad+4(1-2\Delta^{\mu\mu}_{kk})\sqrt{\Delta^{\mu\mu}_{kk}(1-\Delta^{\mu\mu}_{kk})e^{b,\mu\mu}_{y,kk}(1-e^{b,\mu\mu}_{y,kk})},$ where $\alpha,\alpha_{1},\alpha_{2},\beta,\beta_{1},\beta_{2}\in\{0,\nu,\mu\}$, $\gamma=\{x,y\}$ represents either $X$ or $Y$ basis, $p_{k}^{\mu}$, $p_{l}^{\beta}$ are given in Eq. 5, $\epsilon_{\alpha_{1}\alpha_{2}}$ and $\epsilon_{\beta_{1}\beta_{2}}$ come from Eq. 13, and $\Delta^{\mu\mu}_{kk}$ is defined in Eq. 20. Approximate Formulation As above, with the additional constraint that $Y_{\gamma,kl}^{\alpha\beta}=0$, for k,l=7,8,,D1k,l=7,8,...,D-1.
@@ III-B MDI QKD with DPR: Rate analysis
Considering the DPR scenario, the secret key rate for the MDI QKD protocol is lower bounded by [gottesman2004security, cao2020discrete]:
$R^{\text{DPR}}_{\text{MDI}}=\sum_{k=\{0,1\}}Q_{x,kk}^{\mu\mu}[1-h_{2}(e_{x,kk}^{p,\mu\mu})]-fQ_{x}^{\mu\mu}h_{2}(E_{x}^{\mu\mu}),$ (17)
where, in general, $Q_{x,kl}^{\mu\mu}$ and $e_{x,kl}^{p,\mu\mu}$, respectively, are the gain and phase error rate, when Alice and Bob encode $|\lambda_{k}\rangle$ and $|\lambda_{l}\rangle$, both with intensity $\mu$, in the $X$ basis, and $Q_{x}^{\mu\mu}$ and $E_{x}^{\mu\mu}$ are, respectively, the total gain and QBER when intensity $\mu$ is used by both Alice and Bob in the $X$ basis.
@@
where $\gamma=\{x,y\}$ represents $X$ or $Y$ basis, and $Q_{\gamma}^{\alpha\beta}$ and $E_{\gamma}^{\alpha\beta}$ are experimentally observable variables in basis $\gamma$, respectively, representing the total gain and QBER, when Alice and Bob use intensities $\alpha$ and $\beta$. The term $Y_{\gamma,kl}^{\alpha\beta}$ denotes the yield in basis $\gamma$ when Alice and Bob encode the states $|\lambda_{k}\rangle$, with intensity $\alpha$, and $|\lambda_{l}\rangle$, with intensity $\beta$, respectively, and $e_{\gamma,kl}^{b,\alpha\beta}$ is the corresponding bit error rate.
Similar to Eq. 13, in the MDI case, we have the following relationships [cao2015discrete]:
@@
The final step is to bound the phase error rate $e^{p,\mu\mu}_{x,kk}$ in the $X$ basis by the bit error rate $e_{y,kk}^{b,\mu\mu}$ in the $Y$ basis. Similar to what we did in Section III-A, the basis dependence can be characterized by [gottesman2004security, koashi2005simple]:
$\begin{split}\Delta_{kk}^{\mu\mu}=\text{min}(1,\frac{1-F_{k,\text{MDI}}}{2Y_{x,kk}^{\mu\mu}}),\end{split}$ (20)
which corresponds to the case when both Alice and Bob encode $|\lambda_{k}\rangle$. Thereby, in our optimization problem, we only consider the contribution from $e_{x,kk}^{p,\mu}$ for $k=0$ and $k=1$, for which the following upper bound is given [tamaki2003unconditionally, lo2006security, cao2020discrete]:
$\begin{split}e^{p,\mu\mu}_{x,kk}\leq&e^{b,\mu\mu}_{y,kk}+4\Delta^{\mu\mu}_{kk}(1-\Delta^{\mu\mu}_{kk})(1-2e^{b,\mu\mu}_{y,kk})\\ &+4(1-2\Delta^{\mu\mu}_{kk})\sqrt{\Delta^{\mu\mu}_{kk}(1-\Delta^{\mu\mu}_{kk})e^{b,\mu\mu}_{y,kk}(1-e^{b,\mu\mu}_{y,kk})}.\end{split}$ (21)
@@ IV Analytical Bounds
In this section, we present analytical bounds on the key parameters in Eq. 11 and Eq. 17. The key objective is to derive bounds on such parameters using experimentally accessible observables from vacuum, decoy, and signal states. Our approach parallels the methodology used in CPR QKD [ma2005practical, gobby2004quantum, sun2013practical], but requires additional care due to the differences introduced by DPR [cao2020discrete]. Here, we summarize the final results leaving detailed derivations to Appendix A and Appendix B.
@@ IV-A Analytical bounds for BB84 with DPR
In this section, we give analytical bounds on $Y_{\gamma,0}^{\mu}$, $Y_{\gamma,1}^{\mu}$, $W_{y,0}^{\mu}$, and $W_{y,1}^{\mu}$ that contribute to the key rate expression in Eq. 11, $W_{\gamma,k}^{\alpha}=e_{\gamma,k}^{b,\alpha}Y_{\gamma,k}^{\alpha}$.
@@
$W_{y,0}^{\mu}$ $\leq W_{y,0}^{\mu U}\equiv\min(Q_{y}^{0}E_{y}^{0}+\epsilon_{\mu 0},0.5),$ (22)
where $Q_{\gamma}^{0}$ and $Q_{y}^{0}E_{y}^{0}$ are directly obtained from observables. Using these, we can derive an upper bound for the bit error rate as $e_{y,0}^{b,\mu}\leq e_{y,0}^{b,\mu U}\equiv W_{y,0}^{\mu U}/Y_{y,0}^{\mu L}$.
The next key parameters in the analysis are $Y_{\gamma,1}^{\mu}$ and $W_{y,1}^{\mu}=e_{y,1}^{b,\mu}Y_{y,1}^{\mu}$, which correspond to the state $|\lambda_{1}\rangle$. The lower bound on $Y_{\gamma,1}^{\mu}$ is given by:
@@
Next, the upper bound on $W_{y,1}^{\mu}$ is given by
$\begin{split}&W_{y,1}^{\mu}\leq W_{y,1}^{\mu U}\\ &\equiv\frac{e^{\alpha_{1}}Q_{y}^{\alpha_{1}}E_{y}^{\alpha_{1}}-e^{\alpha_{2}}Q_{y}^{\alpha_{2}}E_{y}^{\alpha_{2}}-(e^{\alpha_{1}}p_{0}^{\alpha_{1}}-e^{\alpha_{2}}p_{0}^{\alpha_{2}})Q_{y}^{0}E_{y}^{0}}{e^{\alpha_{1}}p_{1}^{\alpha_{1}}-e^{\alpha_{2}}p_{1}^{\alpha_{2}}}\\ &+\frac{e^{\alpha_{1}}p_{0}^{\alpha_{1}}\epsilon_{\alpha_{1}0}+e^{\alpha_{2}}p_{0}^{\alpha_{2}}\epsilon_{\alpha_{2}0}+e^{\alpha_{1}}p_{1}^{\alpha_{1}}\epsilon_{\mu\alpha_{1}}+e^{\alpha_{2}}p_{1}^{\alpha_{2}}\epsilon_{\mu\alpha_{2}}}{e^{\alpha_{1}}p_{1}^{\alpha_{1}}-e^{\alpha_{2}}p_{1}^{\alpha_{2}}},\end{split}$ (25)
where the condition $e^{\alpha_{1}}p_{1}^{\alpha_{1}}-e^{\alpha_{2}}p_{1}^{\alpha_{2}}>0$ is met when $\alpha_{1}>\alpha_{2}$, which is the case for our choice of $\alpha_{1}=\nu$ and $\alpha_{2}=0$. Consequently, the upper bound of the bit error rate is $e^{b,\mu U}_{y,1}=W_{y,1}^{\mu U}/Y_{y,1}^{\mu L}$.
In summary, by employing the vacuum+weak decoy state method and applying fidelity-based constraints, we are able to derive lower bounds for the yields $Y_{\gamma,0}^{\mu}$ and $Y_{\gamma,1}^{\mu}$, as shown in Eq. 22 and Eq. 23, respectively. Similarly, upper bounds for the bit error rates $e_{y,0}^{b,\mu}$ and $e_{y,1}^{b,\mu}$ are obtained by using Eq. 22 and Eq. 25. With these bounds, the lower bound on the key rate can be calculated by using Eq. 11, where the phase error rate of the $X$ basis is bounded via the bit error rate in the $Y$ basis, as described in Eq. 16. All other required parameters in Eq. 11 are either directly measurable in experiments or computable through simulations.

### v2

@@ III-A BB84 with DPR: Rate analysis
The key rate of the BB84 protocol in Section II-B is lower bounded by [7]
$R^{\text{DPR}}_{\text{BB84}}$ $=\sum_{k=\{0,1\}}Q_{z,k}^{\mu}[1-h_{2}(e_{z,k}^{p,\mu})]-fQ_{z}^{\mu}h_{2}(E_{z}^{\mu}),$ (11)
where $Q_{z,k}^{\mu}$ denotes the gain when Alice encodes the state $|\lambda_{k}\rangle$ with intensity $\mu$, $f$ is the error correction inefficiency, and $h_{2}(x)=-x\log_{2}(x)-(1-x)\log_{2}(1-x)$ is the Shannon binary entropy function. The term $e_{z,k}^{p,\mu}$ denotes the phase error rate for $|\lambda_{k}\rangle$ with intensity $\mu$ as the input state, which in the $Z$ basis corresponds to errors caused by Pauli operator $\sigma_{z}$. The second term in the key rate expression accounts for the cost of error correction, where $Q_{z}^{\mu}$ and $E_{z}^{\mu}$ are the total gain and quantum bit error rate (QBER), respectively, for intensity $\mu$. All above terms are for the $Z$ basis, which is the main basis used in the efficient QKD protocol we defined in Section II-B.
To compute the key rate, two critical parameters must be bounded: The yield term $Y_{z,k}^{\mu}=Q_{z,k}^{\mu}/p_{k}^{\mu}$, and the phase error rate $e^{p,\mu}_{z,k}$. Both parameters are associated with the quantum state $|\lambda_{k}\rangle$, rather than the photon number state $|k\rangle$ in CPR. Another distinct feature in DPR is that, unlike the CPR case, the yield, gain, and error parameters corresponding to $|\lambda_{k}\rangle$ are now functions of the employed intensity as well. This will make the optimization problem more tedious to solve as we see below.
@@
where the subscript $\gamma=\{z,x\}$ specfies the basis and $e_{\gamma,k}^{b,\alpha}$ is the bit error rate when $|\lambda_{k}\rangle$, with intensity $\alpha$, is at the encoder input.
To minimize Eq. 11 subject to Eq. 12, we use the following inequalities [7, 2]:
@@
where $\alpha_{1}$ and $\alpha_{2}$ are two intensity variables in $\{0,\nu,\mu\}$, and the fidelity $F_{\alpha_{1}\alpha_{2}}$ is given by
@@
Finally, to bound the privacy amplification term, the phase error rate term $e^{p,\mu}_{z,k}$ needs to be upper bounded. We use the virtual protocol described in [7, 11] to establish a relationship between the bit error rate $e^{b,\mu}_{x,k}$ in the $X$ basis and the phase error rate $e^{p,\mu}_{z,k}$ in the $Z$ basis, which characterizes the effective basis dependence in realistic scenarios by the following parameter:
$\begin{split}\Delta_{k}^{\mu}=\text{min}(1,\frac{1-F_{k,\text{BB84}}}{2Y_{z,k}^{\mu}}).\end{split}$ (15)
Using a bound proved in [14], we can get a relationship between $e^{b,\mu}_{x,k}$ and $e^{p,\mu}_{z,k}$ as follows:
$e^{p,\mu}_{z,k}$ $\leq e^{b,\mu}_{x,k}+4\Delta^{\mu}_{k}(1-\Delta^{\mu}_{k})(1-2e^{b,\mu}_{x,k})+$
$4(1-2\Delta^{\mu}_{k})\sqrt{\Delta^{\mu}_{k}(1-\Delta^{\mu}_{k})e^{b,\mu}_{x,k}(1-e^{b,\mu}_{x,k})},k=0,1.$ (16)
@@
Box 2: Nonlinear optimization problem for DPR MDI QKD with vacuum+weak decoy states Generic Formulation Minimize $R^{\text{DPR}}_{\text{MDI}}=\sum_{k=\{0,1\}}Q_{z,kk}^{\mu\mu}[1-h_{2}(e_{z,kk}^{p,\mu\mu})]-fQ_{z}^{\mu\mu}h_{2}(E_{z}^{\mu\mu}).$ subject to $Q_{\gamma}^{\alpha\beta}=\sum_{k,l=0}^{D-1}p_{k}^{\alpha}p_{l}^{\beta}Y_{\gamma,kl}^{\alpha\beta},$ $Q_{\gamma}^{\alpha\beta}E_{\gamma}^{\alpha\beta}=\sum_{k,l=0}^{D-1}p_{k}^{\alpha}p_{l}^{\beta}e_{\gamma,kl}^{b,\alpha\beta}Y_{\gamma,kl}^{\alpha\beta},$ $|e_{\gamma,kl}^{b,\alpha_{1}\beta_{1}}Y_{\gamma,kl}^{\alpha_{1}\beta_{1}}-e_{\gamma,kl}^{b,\alpha_{2}\beta_{2}}Y_{\gamma,kl}^{\alpha_{2}\beta_{2}}|\leq\epsilon_{\alpha_{1}\alpha_{2}}+\epsilon_{\beta_{1}\beta_{2}},$ $|Y_{\gamma,kl}^{\alpha_{1}\beta_{1}}-Y_{\gamma,kl}^{\alpha_{2}\beta_{2}}|\leq\epsilon_{\alpha_{1}\alpha_{2}}+\epsilon_{\beta_{1}\beta_{2}},\quad{k,l=0,\ldots,D-1}$ $e^{p,\mu\mu}_{z,kk}\leq e^{b,\mu\mu}_{x,kk}+4\Delta^{\mu\mu}_{kk}(1-\Delta^{\mu\mu}_{kk})(1-2e^{b,\mu\mu}_{x,kk})+$ $4(1-2\Delta^{\mu\mu}_{kk})\sqrt{\Delta^{\mu\mu}_{kk}(1-\Delta^{\mu\mu}_{kk})e^{b,\mu\mu}_{x,kk}(1-e^{b,\mu\mu}_{x,kk})},k=0,1$ where $\alpha,\alpha_{1},\alpha_{2},\beta,\beta_{1},\beta_{2}\in\{0,\nu,\mu\}$, $\gamma=\{z,x\}$ represents either $Z$ or $X$ basis, $p_{k}^{\mu}$, $p_{l}^{\beta}$ are given in Eq. 5, $\epsilon_{\alpha_{1}\alpha_{2}}$ and $\epsilon_{\beta_{1}\beta_{2}}$ come from Eq. 13, and $\Delta^{\mu\mu}_{kk}$ is defined in Eq. 20. Approximate Formulation As above, with the additional constraint that $Y_{\gamma,kl}^{\alpha\beta}=0$, for k,l=7,8,,D1k,l=7,8,...,D-1.
@@ III-B MDI QKD with DPR: Rate analysis
Considering the DPR scenario, the secret key rate for the MDI QKD protocol is lower bounded by [7, 3]:
$R^{\text{DPR}}_{\text{MDI}}=\sum_{k=\{0,1\}}Q_{z,kk}^{\mu\mu}[1-h_{2}(e_{z,kk}^{p,\mu\mu})]-fQ_{z}^{\mu\mu}h_{2}(E_{z}^{\mu\mu}),$ (17)
where, in general, $Q_{z,kl}^{\mu\mu}$ and $e_{z,kl}^{p,\mu\mu}$, respectively, are the gain and phase error rate, when Alice and Bob encode $|\lambda_{k}\rangle$ and $|\lambda_{l}\rangle$, both with intensity $\mu$, in the $Z$ basis, and $Q_{z}^{\mu\mu}$ and $E_{z}^{\mu\mu}$ are, respectively, the total gain and QBER when intensity $\mu$ is used by both Alice and Bob in the $X$ basis. As in the case of BB84, in Eq. 17, we only include terms that correspond to $|\lambda_{0}\rangle$ and $|\lambda_{1}\rangle$. In particular, we have only accounted for the terms that correspond to sending the same state by both Alice and Bob. We can numerically verify that all other terms contribute negligibly to the key rate in regions of interest. The choice of Eq. 17, as the lower bound, also simplifies deriving analytical bounds for the relevant terms.
@@
where $\gamma=\{z,x\}$ represents $Z$ or $X$ basis, and $Q_{\gamma}^{\alpha\beta}$ and $E_{\gamma}^{\alpha\beta}$ are experimentally observable variables in basis $\gamma$, respectively, representing the total gain and QBER, when Alice and Bob use intensities $\alpha$ and $\beta$. The term $Y_{\gamma,kl}^{\alpha\beta}$ denotes the yield in basis $\gamma$ when Alice and Bob encode the states $|\lambda_{k}\rangle$, with intensity $\alpha$, and $|\lambda_{l}\rangle$, with intensity $\beta$, respectively, and $e_{\gamma,kl}^{b,\alpha\beta}$ is the corresponding bit error rate.
Similar to Eq. 13, in the MDI case, we have the following relationships [2]:
@@
The final step is to bound the phase error rate $e^{p,\mu\mu}_{z,kk}$ in the $Z$ basis by the bit error rate $e_{x,kk}^{b,\mu\mu}$ in the $X$ basis. Similar to what we did in Section III-A, the basis dependence can be characterized by [7, 11]:
$\begin{split}\Delta_{kk}^{\mu\mu}=\text{min}(1,\frac{1-F_{k,\text{MDI}}}{2Y_{z,kk}^{\mu\mu}}),\end{split}$ (20)
which corresponds to the case when both Alice and Bob encode $|\lambda_{k}\rangle$. Using the above, it can be shown that [26, 14, 3]:
$\begin{split}&e^{p,\mu\mu}_{z,kk}\leq e^{b,\mu\mu}_{x,kk}+4\Delta^{\mu\mu}_{kk}(1-\Delta^{\mu\mu}_{kk})(1-2e^{b,\mu\mu}_{x,kk})\\ +&4(1-2\Delta^{\mu\mu}_{kk})\sqrt{\Delta^{\mu\mu}_{kk}(1-\Delta^{\mu\mu}_{kk})e^{b,\mu\mu}_{x,kk}(1-e^{b,\mu\mu}_{x,kk})},k=0,1.\end{split}$ (21)
@@ IV Analytical Bounds
In this section, we present analytical bounds on the key parameters in Eq. 11 and Eq. 17. The key objective is to derive bounds on such parameters using experimentally accessible observables from vacuum, decoy, and signal states. Our approach parallels the methodology used in CPR QKD [18, 6, 25], but requires additional care due to the differences introduced by DPR [3]. Here, we summarize the final results leaving detailed derivations to Appendix A and Appendix B.
@@ IV-A Analytical bounds for BB84 with DPR
In this section, we give analytical bounds on $Y_{\gamma,0}^{\mu}$, $Y_{\gamma,1}^{\mu}$, $W_{x,0}^{\mu}$, and $W_{x,1}^{\mu}$, with $W_{\gamma,k}^{\alpha}=e_{\gamma,k}^{b,\alpha}Y_{\gamma,k}^{\alpha}$, which contribute to the key rate expression in Eq. 11.
@@
$W_{yx,0}^{\mu}$ $\leq W_{x,0}^{\mu U}\equiv\min(Q_{x}^{0}E_{x}^{0}+\epsilon_{\mu 0},0.5),$ (22)
where $Q_{\gamma}^{0}$ and $Q_{x}^{0}E_{x}^{0}$ are directly obtained from observables. Using these, we can derive an upper bound for the bit error rate as $e_{x,0}^{b,\mu}\leq e_{x,0}^{b,\mu U}\equiv W_{x,0}^{\mu U}/Y_{x,0}^{\mu L}$.
The next key parameters in the analysis are $Y_{\gamma,1}^{\mu}$ and $W_{x,1}^{\mu}=e_{x,1}^{b,\mu}Y_{x,1}^{\mu}$, which correspond to the state $|\lambda_{1}\rangle$. The lower bound on $Y_{\gamma,1}^{\mu}$ is given by:
@@
Next, the upper bound on $W_{x,1}^{\mu}$ is given by
$\begin{split}&W_{x,1}^{\mu}\leq W_{x,1}^{\mu U}\\ &\equiv\frac{Q_{x}^{\alpha_{1}}E_{x}^{\alpha_{1}}-p_{0}^{\alpha_{1}}Q_{x,0}^{0}E_{x,0}^{0}+p_{0}^{\alpha_{1}}\epsilon_{\alpha_{1}0}+p_{1}^{\alpha_{1}}\epsilon_{\mu\alpha_{1}}}{p_{1}^{\alpha_{1}}}.\end{split}$ (25)
Consequently, the upper bound on the bit error rate is given by $e^{b,\mu U}_{x,1}=W_{x,1}^{\mu U}/Y_{x,1}^{\mu L}$.
In summary, by employing the vacuum+weak decoy state method and applying fidelity-based constraints, we are able to derive lower bounds for the yields $Y_{\gamma,0}^{\mu}$ and $Y_{\gamma,1}^{\mu}$, as shown in Eq. 22 and Eq. 23, respectively. Similarly, upper bounds for the bit error rates $e_{x,0}^{b,\mu}$ and $e_{x,1}^{b,\mu}$ are obtained by using Eq. 22 and Eq. 25. With these bounds, the lower bound on the key rate can be calculated by using Eq. 11, where the phase error rate of the $Z$ basis is bounded via the bit error rate in the $X$ basis, as described in Eq. 16. All other required parameters in Eq. 11 are either directly measurable in experiments or computable through simulations.